# Null-Space Filtering for Data-Free Continual Model Merging: Preserving Transparency, Promoting Fidelity

### v1

@@ 4 Method: NUFILT
Figure 3: Overview of the NUFILT procedure. ❶ Filtering: the new task vector is processed through a null-space projector that suppresses activations from previous tasks, ensuring transparency to past knowledge. ❷ Adapting: within the filter, a lightweight LoRA adapter refines the update for the current task using a data-free objective. ❸ Fusing: the filter, task vector, and LoRA module are merged back into the backbone, keeping the parameter count and inference cost unchanged.
@@ 4.1 Steps in NUFILT: Filtering, Adapting & Fusing
$\tilde{\tau}_{\leq t-1}^{(l)}=\theta_{t-1}^{\text{merged},(l)}-\theta_{0}^{(l)}.$ (7)
@@
which removes components aligned with earlier task directions. Thus, for old task features $h^{(l)}\in\mathrm{span}(\hat{V}_{\leq t-1}^{(l)})$, we have $P_{t}^{(l)}h^{(l)}=0$, ensuring transparency to prior knowledge.
@@ 4.2 Data-free Objective of NUFILT
$\mathcal{L}_{\text{trans}}=\mathbb{E}_{X\sim\mathcal{D}_{i\leq t-1}}\!\|(\tilde{\tau}_{\leq t}-\tilde{\tau}_{\leq t-1})X^{\top}\|_{F}^{2},\quad\mathcal{L}_{\text{fid}}=\mathbb{E}_{X\sim\mathcal{D}_{t}}\!\|(\tilde{\tau}_{\leq t}-\tau_{t})X^{\top}\|_{F}^{2},$ (12)
@@
$T=\begin{bmatrix}\tilde{\tau}_{\leq t-1}^{(l)}\hat{V}_{\leq t-1}^{(l)}\\ \tau_{t}^{(l)}\hat{V}_{t}^{(l)}\end{bmatrix},\quad\hat{V}=\begin{bmatrix}\hat{V}_{\leq t-1}^{(l)}\\ \hat{V}_{t}^{(l)}\end{bmatrix},\quad M=\tilde{\tau}_{\leq t-1}^{(l)}+\tau_{t}^{(l)}P_{t}^{(l)}.$ (17)
Here the residual $\tau_{t}^{(l)}B_{t}^{(l)}A_{t}^{(l)}$ bridges $M\hat{V}$ and $T$, ensuring transparency to past tasks and fidelity to the new one. Since $\tau_{t}^{(l)}$ is generally non-square and low-rank, closed-form for $(A_{t}^{(l)},B_{t}^{(l)})$ are unstable; we thus optimize Eq. 16 via gradient descent. The full procedure is summarized in Algo. 1.
@@
   $A_{t}^{(l)}\leftarrow\mathbf{0},\quad B_{t}^{(l)}\sim\mathcal{N}(0,\sigma_{init}^{2})$ $\triangleright$  initialize LoRA components with low rank $r_{l}$
@@
   compute $M=\tilde{\tau}_{\leq t-1}^{(l)}+\tau_{t}^{(l)}P_{t}^{(l)}$, $T$, and $\hat{V}$ via Eq. 17
   $\mathcal{L}(A_{t}^{(l)},B_{t}^{(l)})=\left\|T-\left(M+\tau_{t}^{(l)}B_{t}^{(l)}A_{t}^{(l)}\right)\hat{V}\right\|_{F}^{2}$ $\triangleright$  data-free objective (Eq. 16)
@@ 5.1 Experimental Setups
Benchmarks and Protocols. We evaluate our approach on both vision and NLP tasks to assess scalability and generality. For vision, following Tang et al. (2025), we adopt CLIP-ViT backbones (Radford et al., 2021) and construct three groups of 8, 14, and 20 image classification tasks. We use publicly available ViT-B/32, ViT-B/16, and ViT-L/14 checkpoints, each fine-tuned on up to 20 datasets (Tang et al., 2024a). To ensure robustness to task order, all experiments are repeated over 10 random permutations (seeds 42–51). For NLP, we evaluate on eight GLUE tasks (Wang et al., 2019) using Flan-T5-base (Chung et al., 2024).
@@
Metrics and Baselines. We evaluate performance with two standard metrics: average accuracy (ACC) and backward transfer (BWT) (Lin et al., 2022). ACC is the mean accuracy of the final merged model across all tasks, $\text{ACC}=\tfrac{1}{T}\sum_{i=1}^{T}a_{i}(\theta_{T}^{\text{merged}}),$ where $a_{i}(\cdot)$ denotes accuracy on task ii [^1] For STSB (NLP), we report Spearman’s $\rho$, denoted by $a_{i}(\cdot)$ for notational consistency.. BWT quantifies the effect of merging on past tasks by comparing their performance before and after the final merge: $\text{BWT}=\tfrac{1}{T-1}\sum_{i=1}^{T-1}\Big[a_{i}(\theta_{T}^{\text{merged}})-a_{i}(\theta_{i}^{\text{merged}})\Big].$ In addition to the methods in Sec. 3.2, we compare against Ties-Merging (Yadav et al., 2023), MagMax-Ind (Marczak et al., 2024), WEMOE (Tang et al., 2024b), and WUDI-Merging (Cheng et al., ). Detailed descriptions are provided in the Appendix B.3.
@@ 5.2 Main Results
Results of Vision Tasks The comparative results on vision tasks are summarized in Tab. 1, highlighting that our method, NUFILT, consistently outperforms prior merging techniques across various model architectures and task sequences, all without needing extra parameters or data access. Relative to baselines like OPCM and WUDI-Merging, NUFILT achieves marked improvements in average accuracy while reducing backward transfer. For example, on ViT-B/32 with 8 tasks, it reaches 83.6% accuracy, outperforming OPCM by 8.1% and WUDI-Merging by 8.9%; for 20 tasks, the gains are 5.3% over OPCM. On ViT-L/14, it surpasses OPCM by 4.6% on 8 tasks and 8.7% on 20 tasks, with backward transfer limited to -1.1% versus -2.6% for OPCM on 8 tasks. Overall, NUFILT closes the gap to the individual fine-tuning benchmark, lagging by only 2.7% on ViT-L/14 for 8 tasks, underscoring its robustness in continual merging settings.
Results of NLP Tasks The results for NLP tasks using the Flan-T5-base model on 8 tasks are detailed in Tab. 2, showing that NUFILT excels all merging strategies. Against comparable baselines such as OPCM and WUDI-Merging, NUFILT delivers notable gains in overall performance. It achieves an average accuracy of 83.7%, exceeding WUDI-Merging by 1.5% and OPCM by 3.1%, while limiting backward transfer to -1.5% compared to -3.9% for WUDI-Merging and -2.5% for OPCM. NUFILT narrows the divide with individual fine-tuning, trailing by merely 2.7%, affirming its efficacy in continual merging for language models.
@@ 5.3 Ablation and Analysis Results
Ablation on Each Component. We conduct an ablation study to disentangle the contributions of the null-space filter and the data-free objective. Results are summarized in Tab. 3. (1) Simply accumulating task vectors without any constraint leads to severe performance degradation. Both accuracy and backward transfer drop sharply as the number of tasks increases, confirming that direct parameter addition suffers from catastrophic interference. (2) Projecting task vectors onto the null-space of previous tasks significantly stabilizes merging. Compared to naive merging, this strategy improves accuracy by more than 20 points across all task counts and reduces forgetting to near zero. (3) Using the data-free objective alone (without null-space filtering) recovers part of the performance but still suffers from negative backward transfer, especially when scaling to more tasks. Combining both components yields the best balance between transparency and fidelity.
(a) Vary $r_{p}$, $r_{l}=r_{v}=0$
(b) Vary $r_{l}$, $r_{p}=128,r_{v}=8$
(c) Vary $r_{v}$, $r_{p}=128,r_{l}=16$
@@
Table 4: Overhead of ViT-B/32 under 8-task continual merging: per-task iterations, total solving time, peak GPU memory, and final parameter count.
Model Iters. Time GPU Mem. Param. ACC(%)
LW AdaMerging 50 47.1s 1.6GB 87.5M 53.4 $\pm$3.2
LoRA-WEMOE 50 72.9s 1.8GB 103.7M 68.8 $\pm$7.8
WUDI-Merging 50 28.9s 2.0GB 87.5M 74.7 $\pm$ 6.6
NUFILT 25 9.3s 1.8GB 87.5M 83.3 $\pm$0.3
100 35.9s 1.8GB 87.5M 83.6 $\pm$0.2
50 18.4s 1.8GB 87.5M 83.6 $\pm$0.2
Adaptation Overhead. Tab. 4 details the adaptation overhead for ViT-B/32 under 8-task continual merging. NUFILT outperforms baselines in both efficiency and accuracy. Compared to LW AdaMerging and LoRA-WEMOE, it uses similar or less GPU memory and runtime while achieving higher accuracy. WUDI-Merging is efficient but lags in accuracy. Remarkably, NUFILT reaches 83.3% accuracy with just 25 iterations and 9.3s, under one-third baseline runtimes. This efficiency underscores NUFILT’s scalability and practicality for continual merging.
@@ 6 Conclusion
In this work, we addressed the problem of data-free continual model merging, where independently fine-tuned models must be consolidated sequentially into a single backbone without access to task data or earlier checkpoints. The central challenge lies in enforcing transparency and fidelity. To this end, we established that task vectors approximately align with representation subspaces, providing a geometric foundation for continual merging. Building on this insight, we proposed NUFILT, a novel framework that integrates null-space filtering to suppress interference with prior knowledge and projection-aware adaptation to recover task-specific fidelity via data-free surrogate objectives. Experiments across vision and NLP benchmarks demonstrated that NUFILT achieves state-of-the-art results, improving over recent methods by 4–7% on average while substantially reducing forgetting and narrowing the gap to individual fine-tuning. We believe this work provides a principled and practical step toward scalable, theoretically grounded solutions for continual model merging.
@@
A   Proof of Theoretical Results ........................................................................................................................................................................14
  A.1  Proof of Theorem 1 ........................................................................................................................................................................14
  A.2  Proof of Corollary 1 ........................................................................................................................................................................16
B   Additional Descriptions ........................................................................................................................................................................17 B.1  Details of Dataset and Task Settings ........................................................................................................................................................................17 B.2  Details of Downstream Models ........................................................................................................................................................................18 B.3  Details of Baselines ........................................................................................................................................................................18 C   Additional Results ........................................................................................................................................................................20 C.1  Detailed Overall Performance Results ........................................................................................................................................................................20 C.2  Extended Visualizations on Subspace Alignment ........................................................................................................................................................................20 D   Discussions ........................................................................................................................................................................25 D.1  Limitations ........................................................................................................................................................................25 D.2  Broader Impacts ........................................................................................................................................................................25 D.3  LLM Usage ........................................................................................................................................................................25
@@ A.1 Proof of Theorem 1
Theorem 1 extends Proposition A.1 (Cheng et al., ) from individual task vectors to subspaces. While Proposition A.1 demonstrates that each task vector can be approximated as a linear combination of representation features, Theorem 1 further establishes approximate alignment between the singular subspaces of task vectors and task representations.
@@
Proposition A.1 (Approximate Linear Combination (Cheng et al., )).
@@
Let $\tau^{(l)}\in\mathbb{R}^{d_{o}\times d_{i}}$ be the task vector at layer $l$, and $H\in\mathbb{R}^{N\times d_{i}}$ a representation matrix of rank $r_{d}$ with right singular vectors $V_{d}\in\mathbb{R}^{d_{i}\times r_{d}}$. Suppose $\tau^{(l)}$ admits the decomposition $\tau^{(l)}=T_{0}+E$ with $T_{0}=BH,\;\mathrm{rank}(T_{0})=r_{d}$, where each row $E_{k}$ of $E$ satisfies $\|E_{k}\|_{2}\leq\Psi_{k}$. If $\sigma_{r_{d}}(T_{0})>\sqrt{d_{o}}\,\max_{k}\Psi_{k},$ then for any $r_{v}\geq r_{d}$, the span of the top $r_{v}$ right singular vectors $\hat{V}$ of $\tau^{(l)}$ is approximately aligned with $\mathrm{span}(V_{d})$ in the sense that $\mathcal{A}(V_{d}^{(l)},\hat{V}^{(l)})\;\leq\;\zeta^{2},$ (21) where $\mathcal{A}(V_{d}^{(l)},\hat{V}^{(l)})=1-\tfrac{1}{r_{d}}\|\hat{V}^{\top}V_{d}\|_{F}^{2}$ and $\zeta^{2}=\left(\frac{\sqrt{d_{o}}\,\max_{k}\Psi_{k}}{\sigma_{r_{d}}(T_{0})-\sqrt{d_{o}}\,\max_{k}\Psi_{k}}\right)^{2}.$
@@
[…]
@@
$\mathcal{A}(V_{d},\hat{V}_{r_{d}})=\sqrt{\frac{1}{r_{d}}\sum_{j=1}^{r_{d}}\cos^{2}\theta_{j}}\;\geq\;\cos\theta_{\max}=\sqrt{1-\sin^{2}\theta_{\max}}.$ (25)
Using $1-\sqrt{1-u}\leq u$ for $u\in[0,1]$, we obtain
$1-\mathcal{A}(V_{d},\tilde{V}_{r_{d}})\leq\sin^{2}\theta_{\max}\leq\left(\frac{\sqrt{d_{o}}\,\max_{k}\Psi_{k}}{\sigma_{r_{d}}(T_{0})-\sqrt{d_{o}}\,\max_{k}\Psi_{k}}\right)^{2}.$ (26)
Finally, since enlarging the subspace only increases the overlap, for any $r_{v}\geq r_{d}$,
$\mathcal{A}(V_{d},\hat{V})\geq\mathcal{A}(V_{d},\hat{V}_{r_{d}}),$ (27)
so the same bound holds. ∎

### v2

@@ 4 Method: NUFILT
Figure 3: Overview of the NUFILT procedure. ❶ Filtering: the new task vector is processed through a null-space projector that suppresses activations from previous tasks, ensuring stability to past knowledge. ❷ Adapting: within the filter, a lightweight LoRA adapter refines the update for the current task using a data-free objective. ❸ Fusing: the filter, task vector, and LoRA module are merged back into the backbone, keeping the parameter count and inference cost unchanged.
@@ 4.1 Steps in NUFILT: Filtering, Adapting & Fusing
$\tilde{\tau}_{\leq t-1}^{(l)}=\theta_{t-1}^{\text{merged},(l)}-\theta_{0}^{(l)},$ (7)
@@
which removes components aligned with earlier task directions. Thus, for old task features $x^{(l)}\in\mathrm{span}(\hat{V}_{\leq t-1}^{(l)})$, we have $P_{t}^{(l)}x^{(l)}=0$, ensuring stability to prior knowledge.
@@ 4.2 Data-free Objective of NUFILT
$\mathcal{L}_{\text{stab}}=\mathbb{E}_{X\sim\mathcal{D}_{i\leq t-1}}\!\|(\tilde{\tau}_{\leq t}-\tilde{\tau}_{\leq t-1})X^{\top}\|_{F}^{2},\quad\mathcal{L}_{\text{plas}}=\mathbb{E}_{X\sim\mathcal{D}_{t}}\!\|(\tilde{\tau}_{\leq t}-\tau_{t})X^{\top}\|_{F}^{2},$ (12)
@@
$\mathcal{T}=\begin{bmatrix}\tilde{\tau}_{\leq t-1}^{(l)}\hat{V}_{\leq t-1}^{(l)}\\ \tau_{t}^{(l)}\hat{V}_{t}^{(l)}\end{bmatrix},\quad\hat{V}=\begin{bmatrix}\hat{V}_{\leq t-1}^{(l)}\\ \hat{V}_{t}^{(l)}\end{bmatrix},\quad M=\tilde{\tau}_{\leq t-1}^{(l)}+\tau_{t}^{(l)}P_{t}^{(l)}.$ (17)
Here the residual $\tau_{t}^{(l)}B_{t}^{(l)}A_{t}^{(l)}$ bridges $M\hat{V}$ and $\mathcal{T}$, ensuring stability to past tasks and plasticity to the new one. Since $\tau_{t}^{(l)}$ is generally non-square and low-rank, closed-form for $(A_{t}^{(l)},B_{t}^{(l)})$ are unstable; we thus optimize Eq. 16 via gradient descent. The full procedure is summarized in Algo. 1.
@@
   $A_{t}^{(l)}\sim\mathcal{N}(0,\sigma_{init}^{2}),\quad B_{t}^{(l)}\leftarrow\mathbf{0}$ $\triangleright$  initialize LoRA components with low rank $r_{l}$
@@
   compute $M=\tilde{\tau}_{\leq t-1}^{(l)}+\tau_{t}^{(l)}P_{t}^{(l)}$, $\mathcal{T}$, and $\hat{V}$ via Eq. 17
   $\mathcal{L}(A_{t}^{(l)},B_{t}^{(l)})=\left\|\mathcal{T}-\left(M+\tau_{t}^{(l)}B_{t}^{(l)}A_{t}^{(l)}\right)\hat{V}\right\|_{F}^{2}$ $\triangleright$  data-free objective (Eq. 16)
@@ 5.1 Experimental Setups
Benchmarks and Protocols. We evaluate our approach on both vision, NLP tasks, and multimodal tasks to assess scalability and generality. For vision, following (Ilharco et al., 2023; Wang et al., ), we adopt CLIP-ViT backbones (Radford et al., 2021) and construct three groups of 8, 14, and 20 image classification tasks. We use publicly available ViT-B/32, ViT-B/16, and ViT-L/14 checkpoints, each fine-tuned on up to 20 datasets (Tang et al., 2024a). To ensure robustness to task order, all experiments are repeated over 10 random permutations (seeds 42–51). For NLP, we evaluate on eight GLUE tasks (Wang et al., 2019) using Flan-T5-base (Chung et al., 2024). For multimodal tasks, we evaluate on LLaVA-1.5-7B (Liu et al., 2023), a widely used vision–language model for instruction following. We adopt four multimodal generative benchmarks: IconQA (Lu et al., 2021), ImageNet captioning (Deng et al., 2009), ScienceQA (Lu et al., 2022), and VizWiz-Caption (Gurari et al., 2018), covering diagram reasoning, natural images, scientific QA, and low-vision real-world scenarios.
@@
Metrics and Baselines. We evaluate performance with two standard metrics: average accuracy (ACC) and backward transfer (BWT) (Lin et al., 2022). ACC is the mean accuracy of the final merged model across all tasks, $\text{ACC}=\tfrac{1}{T}\sum_{i=1}^{T}a_{i}(\theta_{T}^{\text{merged}}),$ where $a_{i}(\cdot)$ denotes accuracy on task ii [^1] For STSB (NLP), we report Spearman’s $\rho$, denoted by $a_{i}(\cdot)$ for notational consistency.. BWT quantifies the effect of merging on past tasks by comparing their performance before and after the final merge: $\text{BWT}=\tfrac{1}{T-1}\sum_{i=1}^{T-1}\Big[a_{i}(\theta_{T}^{\text{merged}})-a_{i}(\theta_{i}^{\text{merged}})\Big].$ In addition to the methods in Sec. 3.2, we compare against Ties-Merging (Yadav et al., 2023), MagMax-Ind (Marczak et al., 2024), WEMOE (Tang et al., 2024b), and WUDI-Merging (Cheng et al., 2025). Detailed descriptions are provided in the Appendix B.3.
@@ 5.2 Main Results
Results on Vision Tasks. The comparative results on vision tasks are summarized in Tab. 1, highlighting that our method, NUFILT, consistently outperforms prior merging techniques across various model architectures and task sequences, all without needing extra parameters or data access. Relative to baselines like OPCM and WUDI-Merging, NUFILT achieves marked improvements in average accuracy while reducing backward transfer. For example, on ViT-B/32 with 8 tasks, it reaches 83.6% accuracy, outperforming OPCM by 8.1% and WUDI-Merging by 8.9%; for 20 tasks, the gains are 5.3% over OPCM. On ViT-L/14, it surpasses OPCM by 4.6% on 8 tasks and 8.7% on 20 tasks, with backward transfer limited to -1.1% versus -2.6% for OPCM on 8 tasks. Overall, NUFILT closes the gap to the individual fine-tuning benchmark, lagging by only 2.7% on ViT-L/14 for 8 tasks, underscoring its robustness in continual merging settings.
Results on NLP Tasks. The results for NLP tasks using the Flan-T5-base model on 8 tasks are detailed in Tab. 2, showing that NUFILT excels all merging strategies. Against comparable baselines such as OPCM and WUDI-Merging, NUFILT delivers notable gains in overall performance. It achieves an average accuracy of 83.7%, exceeding WUDI-Merging by 1.5% and OPCM by 3.1%, while limiting backward transfer to -1.5% compared to -3.9% for WUDI-Merging and -2.5% for OPCM. NUFILT narrows the divide with individual fine-tuning, trailing by merely 2.7%, affirming its efficacy in continual merging for language models.
Results on Multimodal tasks. The results for multimodal tasks using LLaVA-1.5-7B on four benchmarks are summarized in Tab. 3. Compared with strong baselines, NUFILT attains clear improvements in average performance. It reaches an overall score of 70.5%, surpassing OPCM by 2.9%, underscoring its effectiveness in continually merging multimodal models.
@@
Table 3: Results of continual merging LLaVA models on 4 tasks, ordered alphabetically.
Method IconQA Image SciQA VizWiz Average
Individual 75.5 96.0 83.7 64.8 80.0
Pre-trained 14.1 40.9 61.7 41.0 39.4
Task Arithmetic 47.6 67.9 74.2 46.6 59.1
Ties-Merging 69.6 74.4 75.5 44.6 66.0
OPCM 58.0 86.4 81.1 44.9 67.6
NUFILT (Ours) 59.6 93.6 80.6 48.3 70.5
@@ 5.3 Ablation and Analysis Results
Ablation on Each Component. We conduct an ablation study to disentangle the contributions of the null-space filter and the data-free objective. Results are summarized in Tab. 4. (1) Simply accumulating task vectors without any constraint leads to severe performance degradation. Both accuracy and backward transfer drop sharply as the number of tasks increases, confirming that direct parameter addition suffers from catastrophic interference. (2) Projecting task vectors onto the null-space of previous tasks significantly stabilizes merging. Compared to naive merging, this strategy improves accuracy by more than 20 points across all task counts and reduces forgetting to near zero. (3) Using the data-free objective alone (without null-space filtering) recovers part of the performance but still suffers from negative backward transfer, especially when scaling to more tasks. Combining both components yields the best balance between stability and plasticity.
@@
Evaluation of stability and plasticity. Following Eqs. 3 and 4, we compute $\mathcal{L}^{(l)}_{\text{stab}}$ and $\mathcal{L}^{(l)}_{\text{plas}}$ layer-wise and report their averages over all layers. $\mathcal{L}^{(l)}_{\text{stab}}$ measures activation changes on earlier tasks, and $\mathcal{L}^{(l)}_{\text{plas}}$ measures deviation from the individual model of the current task $t$. Since both are losses, lower values indicate better stability and plasticity. As shown in Fig. 5, NUFILT achieves much lower stability loss, indicating minimal disturbance to earlier tasks, and consistently lower plasticity loss, showing closer alignment with the individual solution. The harmonic mean further confirms that NUFILT achieves the best balance between the two criteria.
Figure 5: Evaluation of stability and plasticity surrogates, and their harmonic mean on the CLIP ViT-B/32 8-task continual merging protocol. Lower values in all metrics indicate better performance.
Table 5: Overhead of ViT-B/32 under 8-task continual merging: per-task iterations, solving time, total time, peak GPU memory, and final parameter count.
Model Iters. Solv. Tot. GPU Mem. Param. ACC(%)
TA - - 8.3s - 87.5M 67.5 $\pm$ 0.0
OPCM - - 76.6s 1.3GB 87.5M 75.5 $\pm$ 0.5
LW AdaMerging 50 47.1s 724.2s 1.6GB 87.5M 53.4 $\pm$3.2
LoRA-WEMOE 50 72.9s 754.6s 1.8GB 103.7M 68.8 $\pm$7.8
WUDI-Merging 50 28.9s 37.7s 2.0GB 87.5M 74.7 $\pm$ 6.6
NUFILT 25 9.3s 129.8s 1.8GB 87.5M 83.3 $\pm$0.3
100 35.9s 156.4s 1.8GB 87.5M 83.6 $\pm$0.2
50 18.4s 138.9s 1.8GB 87.5M 83.6 $\pm$0.2
Computation Overhead. Tab. 5 compares the computation overhead of three categories of methods. Training-free approaches (TA, OPCM) are the most efficient since they require no optimization or data loading, but their accuracy is limited. Test-time methods (LW AdaMerging, LoRA-WEMOE) exhibit the largest overall runtime because both data loading and multi-step training add substantial latency. Compared with WUDI-Merging, NUFILT achieves lower solving time because its LoRA-based updates operate on fewer parameters, though its total time is slightly higher owing to the additional SVD computation. Overall, NUFILT provides the best accuracy while maintaining competitive efficiency.
@@ 6 Conclusion
In this work, we addressed the problem of data-free continual model merging, where independently fine-tuned models must be consolidated sequentially into a single backbone without access to task data or earlier checkpoints. The central challenge lies in enforcing stability and plasticity. To this end, we established that task vectors approximately align with representation subspaces, providing a geometric foundation for continual merging. Building on this insight, we proposed NUFILT, a novel framework that integrates null-space filtering to suppress interference with prior knowledge and projection-aware adaptation to recover task-specific plasticity via data-free surrogate objectives. Experiments across vision and NLP benchmarks demonstrated that NUFILT achieves state-of-the-art results, improving over recent methods by 4–7% on average while substantially reducing forgetting and narrowing the gap to individual fine-tuning. We believe this work provides a principled and practical step toward scalable, theoretically grounded solutions for continual model merging.
Acknowledgments
This work was supported in part by National Science and Technology Major Project (2021ZD0112001), National Natural Science Foundation of China (No.62271119, 08120002, U23A20286, 62071086), the Key Research and Development Project of Hainan Province (Grant No. ZDYF2024(LALH)003), Liaoning Province Grant XLYC2403174, the Natural Science Foundation of Sichuan Province under Grant 2025ZNSFSC0475.
@@ Acknowledgments
A   Proof of Theoretical Results ........................................................................................................................................................................16
  A.1  Proof of Theorem 1 ........................................................................................................................................................................16
  A.2  Proof of Corollary 1 ........................................................................................................................................................................18
B   Additional Descriptions ........................................................................................................................................................................19 B.1  Details of Dataset and Task Settings ........................................................................................................................................................................19 B.2  Details of Downstream Models ........................................................................................................................................................................20 B.3  Details of Baselines ........................................................................................................................................................................21 B.4  Hyper-parameter Adjustment Guidelines ........................................................................................................................................................................22 C   Additional Results ........................................................................................................................................................................22 C.1  Detailed Overall Performance Results ........................................................................................................................................................................23 C.2  Extended Visualizations on Subspace Alignment ........................................................................................................................................................................23 D   Discussions ........................................................................................................................................................................23 D.1  Limitations ........................................................................................................................................................................23 D.2  Broader Impacts ........................................................................................................................................................................23 D.3  LLM Usage ........................................................................................................................................................................23
@@ A.1 Proof of Theorem 1
Theorem 1 extends Proposition A.1 (Cheng et al., 2025) from individual task vectors to subspaces. While Proposition A.1 demonstrates that each task vector can be approximated as a linear combination of representation features, Theorem 1 further establishes approximate alignment between the singular subspaces of task vectors and task representations.
@@
Proposition A.1 (Approximate Linear Combination (Cheng et al., 2025)).
@@
Let $\tau^{(l)}\in\mathbb{R}^{d_{o}\times d_{i}}$ be the task vector at layer $l$, and $H\in\mathbb{R}^{N\times d_{i}}$ a representation matrix of rank $r_{d}$ with right singular vectors $V_{d}\in\mathbb{R}^{d_{i}\times r_{d}}$. Suppose $\tau^{(l)}$ admits the decomposition $\tau^{(l)}=T_{0}+E$ with $T_{0}=BH,\;\mathrm{rank}(T_{0})=r_{d}$, where each row $E_{k}$ of $E$ satisfies $\|E_{k}\|_{2}\leq\Psi_{k}$. Let $\sigma_{r_{d}}(T_{0})$ denote the $r_{d}$-th largest singular value of the matrix $T_{0}$. If $\sigma_{r_{d}}(T_{0})>\sqrt{d_{o}}\,\max_{k}\Psi_{k},$ then for any $r_{v}\geq r_{d}$, the span of the top $r_{v}$ right singular vectors $\hat{V}$ of $\tau^{(l)}$ is approximately aligned with $\mathrm{span}(V_{d})$ in the sense that $1-\mathcal{A}(V_{d}^{(l)},\hat{V}^{(l)})\;\leq\;\zeta^{2},$ (21) where $\mathcal{A}(V_{d}^{(l)},\hat{V}^{(l)})=\tfrac{1}{r_{d}}\|\hat{V}^{\top}V_{d}\|_{F}^{2}$ and $\zeta=\frac{\sqrt{d_{o}}\,\max_{k}\Psi_{k}}{\sigma_{r_{d}}(T_{0})-\sqrt{d_{o}}\,\max_{k}\Psi_{k}}.$
@@
(1) Exact case ($E=0$). In this case, $\tau^{(l)}=T_{0}=BH$. Since $H$ has rank $r_{d}$, the right singular space of $H$ is exactly $\mathrm{span}(V_{d})$. Under the rank assumption $\mathrm{rank}(T_{0})=r_{d}$ (which holds when $d_{o}\geq r_{d}$ and $B$ has sufficient rank, as motivated by the NTK regime), $T_{0}$ and $H$ share the same $r_{d}$-dimensional right singular subspace. Hence $\mathrm{span}(\hat{V}_{r_{d}})=\mathrm{span}(V_{d})$, so for any $r_{v}\geq r_{d}$, $\mathcal{A}(V_{d},\hat{V})=1$.
@@
$\mathcal{A}(V_{d},\hat{V}_{r_{d}})=\frac{1}{r_{d}}\sum_{j=1}^{r_{d}}\cos^{2}\theta_{j}\;\geq\;1-\sin^{2}\theta_{\max}.$ (25)
Thus, we obtain
$1-\mathcal{A}(V_{d},\hat{V}_{r_{d}})\;\leq\;\sin^{2}\theta_{\max}\;\leq\;\left(\frac{\sqrt{d_{o}}\,\max_{k}\Psi_{k}}{\sigma_{r_{d}}(T_{0})-\sqrt{d_{o}}\,\max_{k}\Psi_{k}}\right)^{2}=\zeta^{2}.$ (26)
Finally, since enlarging the subspace only increases the overlap,
$\mathcal{A}(V_{d},\hat{V})\geq\mathcal{A}(V_{d},\hat{V}_{r_{d}}).$ (27)
Hence, for any $r_{v}\geq r_{d}$,
[…]